OpenStudy (anonymous):

Multiply the rational number. (x+2/2x)*(x+3/x-1) Explain how to multiply your rational expression.

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

\[\frac{ x+2 }{ 2x }\times \frac{ x+3 }{ x-1 }\] Is that the right problem?

OpenStudy (anonymous):

Yes it is!

OpenStudy (anonymous):

When multiplying fractions, you multiply numerator x numerator and denominator x denominator. You'll have to use the distributive property to simplify, but the new problem is: \[\frac{(x+2)(x+3)}{2x(x-1)}\]

OpenStudy (anonymous):

Sorry I haven't replied in so long @jwhite12 but is this the correct answer?